Anne has $2.35 in nickels and dimes, 27 coins in all. how many nickels does she have? put as an equation
Answer by Earlsdon(6294)

If N = the number of nickels ($0.05) and D = the number of dimes ($0.10), then:
1) N+D = 27 "Anne has...27 coins in all."
2) ($0.05)N+($0.10)D = $2.35 "Anne has $2.35 in nickels and dimes,..."
Rewrite equation 1) as:
1a) N = 27-D and substitute this for N in equation 2).
2a) ($0.05)(27-D)+($0.10)D = $2.35 Simplify and solve for D.
2b) 1.35-0.05D+0.1D = 2.35
2c) 0.05D = 1.00
2d) D = 20 and N = 27-20 = 7
Anne has 7 nickels.